the length of a rectangle is 6 meters less than 4 times the width. if the perimeter is 38 meters, find the length and the width of the rectangie
Answer by math_helper(2461)


L = 4W - 6

2L + 2W = 38
2(4W-6) + 2W = 38
10W - 12 = 38
W = 5 --> L = 4(5)-6 = 14

Ans: L=14m, W=5m


Check: (Perimeter) 2(14) + 2(5) = 28+10 = 38